A)Diffraction also occurs with sound waves. Consider 1300 Hz sound waves diffracted by a door that is 85cm wide. What is the angle between the two first-order diffraction minima?

B)What width single slit will produce first-order diffraction minima at angles of ±21 ∘ from the central maximum with 690 nm light?
Answer #1

A) sin(theta) = lamda/a

wavelength lamda = v/f = 343/(1300) = 0.263 m

sin(tehta) = lamda/a = 0.263/0.85 =0.309

theta = sin^(-1)(0.309) =18 degrees


so required angle is 2*theta = 2*18 = 36 degrees

-----------------------------------------

B) width of slit is a = lamda/sin(theta) = 690*10^-9/sin(21) =1.92*10^-6 m = 1.92 micro m